OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = 6, an = 2 • an-1 6, 12, 14, 16, 18, 20 12, 24, 48, 96, 192, 384 0, 2, 12, 14, 16, 18 6, 12, 24, 48, 96, 192

OpenStudy (anonymous):

its the same as the previous one, try it n = 2 what do you get?

OpenStudy (anonymous):

im so confused... @surdawi

OpenStudy (anonymous):

\[a_{n}=2*a_{n-1}\] plug n = 2 and write it down then i will help ypu

OpenStudy (anonymous):

a2=2*a2-1 @surdawi

OpenStudy (anonymous):

2-1=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

so is it \[a_{1}\]?

OpenStudy (anonymous):

a2=2*a1-1 ?? @surdawi

OpenStudy (anonymous):

no i mean what is the value of it

OpenStudy (anonymous):

of a1? @surdawi

OpenStudy (anonymous):

1

OpenStudy (anonymous):

look at the question

OpenStudy (anonymous):

your confusing me O.o @surdawi

OpenStudy (anonymous):

a1 = 6 plug it in

OpenStudy (anonymous):

a2=2*6-1

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

6-1? \[a_{1}=6\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

2*6

OpenStudy (anonymous):

12

OpenStudy (anonymous):

now plug in n = 3